# Investing is Compression

### 14 Apr 2026 v2

@@ Abstract.
In 1956 John Kelly wrote a paper at Bell Labs describing the relationship between gambling and Information Theory. What became known as the Kelly criterion is an objective or utility function and a closed form solution in simple cases. The economist Paul Samuelson argued that it was an arbitrary utility function, and he successfully kept it out of mainstream economics. But he was wrong. We now know, largely through the work of Tom Cover at Stanford, that Kelly’s proposal is objectively optimal: it maximizes long-term wealth, it minimizes the risk of ruin, and in a game-theoretic sense, it is competitively optimal, even over the short term.
One of Cover’s most surprising contributions to portfolio theory was the universal portfolio, related to universal compression in information theory, which performs asymptotically as well as the best constant-rebalanced portfolio in hindsight. Although the algorithm itself is very abstract, one of the key steps Cover used – rewriting the multi-period investing problem as a sum of products rather than a product of sums – reveals the information structure of the investing problem, making it accessible to the techniques of information theory.
That same technique is applied here to show that even in the most general form, Kelly’s objective factors the investing problem into three terms: a money term, an entropy term, and a divergence term. Because the first two terms are independent of the allocation, the only way to maximize the compounding growth rate is to minimize the friction from the divergence term, which measures, in bits, the difference between the chosen distribution and the unknown true distribution. This means that investing is, fundamentally, a compression problem.
This decomposition also yields new practical results. Market data is notoriously non-stationary, and any backtest average is sensitive to extreme and atypical historical events, making it difficult to interpret in a statistical sense. This is true even of log returns. But because the money and entropy terms are constant across strategies in a given backtest, the difference in log growth between two strategies measures, in bits, the change in divergence relative to some unknown ideal. I also introduce a winner fraction heuristic which allocates capital in proportion to each asset’s probability of dominating the candidate set. The growth shortfall of this heuristic relative to the optimal portfolio is bounded by the entropy of the winner fraction distribution, making it especially effective when the candidate set is small or the returns are extreme. To my knowledge, both the heuristic and the entropy bound are original contributions.
@@ 5.3. Winner Fraction and VC-style Investing
(64) $\frac{R_{n}(W^{*})}{R_{n}(W^{\prime})}=\frac{e^{-n\operatorname{H}(W)}\sum_{s\in\mathcal{T}_{n}(W)}r_{s}}{e^{-n\operatorname{H}(W^{\prime})}\sum_{s\in\mathcal{T}_{n}(W^{\prime})}r_{s}}$
@@
(66) $\frac{R_{n}(W^{*})}{R_{n}(W^{\prime})}$ $\leq\frac{e^{-n\operatorname{H}(W)}\sum_{s\in\mathcal{T}_{n}(W)}r_{s}}{e^{-n\operatorname{H}(W^{\prime})}r_{\text{max}}}$

### 15 Apr 2026 v3

@@ Abstract.
In 1956 John Kelly wrote a paper at Bell Labs describing the relationship between gambling and Information Theory. What came to be known as the Kelly Criterion is both an objective and a closed-form solution to sizing wagers when odds and edge are known. Samuelson argued it was arbitrary and subjective, and successfully kept it out of mainstream economics. Luckily it lived on in computer science, mostly because of Tom Cover’s work at Stanford. He showed that it is the uniquely optimal way to invest: it maximizes long-term wealth, minimizes the risk of ruin, and is competitively optimal in a game-theoretic sense, even over the short term.
One of Cover’s most surprising contributions to portfolio theory was the universal portfolio. Related to universal compression in information theory, it performs asymptotically as well as the best constant-rebalanced portfolio in hindsight. I borrow a trick from that algorithm to show that Kelly’s objective, even in the general form, factors the investing problem into three terms: a money term, an entropy term, and a divergence term. The only way to maximize growth is to minimize divergence which measures the difference between our distribution and the true distribution in bits. Investing is, fundamentally, a compression problem.
This decomposition also yields new practical results. Because the money and entropy terms are constant across strategies in a given backtest, the difference in log growth between two strategies measures their relative divergence in bits. I also introduce a winner fraction heuristic which allocates capital in proportion to each asset’s probability of dominating the candidate set. The growth shortfall of this heuristic relative to the optimal portfolio is bounded by the entropy of the winner fraction distribution. To my knowledge, both the heuristic and the entropy bound are original contributions.
@@ 5.3. Winner Fraction and VC-style Investing
(64) $\frac{R_{n}(W^{*})}{R_{n}(W^{\prime})}=\frac{e^{-n\operatorname{H}(W^{*})}\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}}{e^{-n\operatorname{H}(W^{\prime})}\sum_{s\in\mathcal{T}_{n}(W^{\prime})}r_{s}}$
@@
(66) $\frac{R_{n}(W^{*})}{R_{n}(W^{\prime})}$ $\leq\frac{e^{-n\operatorname{H}(W^{*})}\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}}{e^{-n\operatorname{H}(W^{\prime})}r_{\text{max}}}$